a given mass of gas occupies a volume of 950cm³ at 420/c and 0.92*10⁵ pressure. calculate the volume of the gas at S. T. P

1
Expert's answer
2022-02-14T17:07:09-0500

P1 =  0.92×105 N m-2 = 0.92×10Pa = 92000 Pa

T1 = 420 K

V1 = 950 cm3

STP = Standard Temperature and Pressure (T = 273.15 K and P = 101325 Pa)

Therefore,

P2 = 101325 Pa (at STP)

T2 = 273.15 K (at STP)

V2 = unknown (at STP)


Solution:

The Combined Gas Law can be used:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (92000 Pa × 950 cm3 × 273.15 K) / (101325 Pa × 420 K) = 560.98 cm3 = 561 cm3

V2 = 561 cm3


Answer: The volume of the gas at STP is 561 cm3